### v1

- SGM
- score-based generative model
- SNR
- signal-to-noise ratio
- GAN
- generative adversarial network
- VAE
- variational autoencoder
- DDPM
- denoising diffusion probabilistic model
- STFT
- short-time Fourier transform
- iSTFT
- inverse short-time Fourier transform
- SDE
- stochastic differential equation
- ODE
- ordinary differential equation
- OU
- Ornstein-Uhlenbeck
- VE
- Variance Exploding
- DNN
- deep neural network
- PESQ
- Perceptual Evaluation of Speech Quality
- SE
- speech enhancement
- T-F
- time-frequency
- ELBO
- evidence lower bound
- RIR
- room impulse response
- SNR
- signal-to-noise ratio
- LSTM
- long short-term memory
- POLQA
- Perceptual Objectve Listening Quality Analysis
- SDR
- signal-to-distortion ratio
- LSD
- log-spectral distance
- SI-SDR
- scale invariant signal-to-distortion ratio
- ESTOI
- Extended Short-Term Objective Intelligibility
- TCN
- temporal convolutional network
- DRR
- direct-to-reverberant ratio
- NFE
- number of function evaluations
- RTF
- real-time factor
- MOS
- mean opinion scores
- EMA
- expoential moving average
- CNF
- continuous normalizing flow
- ViSQOL
- Virtual Speech Quality Objective Listener
- SRCC
- Spearman’s rank correlation coefficient
- PCC
- Pearson correlation coefficient
- VB-DMD
- VoiceBank-DEMAND
- SNR
- signal-to-noise ratio
Non-intrusive Speech Quality Assessment with Diffusion Models Trained on Clean Speech ††thanks: We acknowledge the support by the German Research Foundation (DFG) in the transregio project Crossmodal Learning (TRR 169) and DASHH (Data Science in Hamburg – Helmholtz Graduate School for the Structure of Matter) with Grant-No. HIDSS-0002.
Abstract
Diffusion models have found great success in generating high quality, natural samples of speech, but their potential for density estimation for speech has so far remained largely unexplored. In this work, we leverage an unconditional diffusion model trained only on clean speech for the assessment of speech quality. We show that the quality of a speech utterance can be assessed by estimating the likelihood of a corresponding sample in the terminating Gaussian distribution, obtained via a deterministic noising process. The resulting method is purely unsupervised, trained only on clean speech, and therefore does not rely on annotations. Our diffusion-based approach leverages clean speech priors to assess quality based on how the input relates to the learned distribution of clean data. Our proposed log-likelihoods show promising results, correlating well with intrusive speech quality metrics such as POLQA and SI-SDR.
Index Terms:
Speech quality assessment, diffusion modelsI Introduction
Speech quality estimation is paramount for evaluating algorithms that tackle speech processing tasks, such as speech enhancement, coding, and synthesis. The golden standard for speech quality estimation is widely considered as the mean opinion scores (MOS) obtained during listening experiments, where participants are asked to rate audio samples. However, listening experiments are expensive, time-consuming and can suffer from listener bias if the instructions are not adequately designed. For these reasons, many instrumental metrics have been proposed to attempt to mimic the result of such listening experiments.
Instrumental metrics can be handcrafted, which include signal-based metrics such as scale invariant signal-to-distortion ratio (SI-SDR)[1] or signal-to-noise ratio (SNR), as well as perceptual metrics integrating some modeling of the human auditory model, like Perceptual Evaluation of Speech Quality (PESQ) [2], its successor Perceptual Objectve Listening Quality Analysis (POLQA)[3] or Virtual Speech Quality Objective Listener (ViSQOL)[4]. Typically, these metrics are intrusive, i.e. they require a reference clean speech signal matching the test utterance. Recording such a clean reference is, however, impractical in real-life scenarios.
In order to avoid relying on reference speech at inference, learning-based metrics based on deep neural networks have been proposed, requiring reference speech only at training. These methods typically try to predict the MOS provided in large labeled speech datasets. These include for example DNSMOS [5], NISQA[6] and NORESQA-MOS [7]. However these do not completely solve the issue of predicting the quality of speech in the wild, as the types of corruption can be unseen during training. Furthermore, supervised methods require large labeled datasets, which can be either inaccessible to the speech research community, or are susceptible to low-quality annotations.
In this paper, we focus on predicting speech quality in an unsupervised fashion, i.e. training our method only on clean speech data. Similar works include SpeechLMScore [8] and VQScore [9]. SpeechLMScore leverages a language model trained on clean speech tokens and computes the likelihood of the test speech sequence according to the language model vocabulary [8]. Lower likelihood will then indicate that the test speech deviates from the clean speech representation of the language model, thereby suggesting low speech quality. In VQScore [7], the authors suggest to train a vector-quantized variational autoencoder (VAE) on clean speech, and inspect the quantization error at the bottleneck of the model. Since the quantized units define a coarse representation of clean speech, observing a large quantization error suggests that the input speech is not well represented by the discrete codebook and therefore it should be considered of low quality.
In this work, we follow similar ideas as SpeechLMScore and VQScore, but instead choose to use a diffusion model [10, 11] for providing us with a representation of clean speech. We compute the likelihood of a test speech by integrating a specific ordinary differential equation (ODE) which, as Song et al. showed [11], provides an exact computation of likelihood using a trained diffusion model. Given that the diffusion model was only trained on clean speech, a test speech utterance of low quality will map to a low-likelihood sample in the terminating Gaussian distribution. In a recent work published during the preparation of this manuscript, Emura [12] showed that the variance of multiple clean speech estimates produced by a diffusion model can be used successfully to estimate the output SI-SDR. However, this work only tests the method on clean speech estimates produced by diffusion-based approaches with the same backbone architecture as in the diffusion models that produce the scores. In comparison to SpeechLMScore [8], our method has no dependence to a choice of speech tokenizer. Rather than using a compressed VAE latent as in VQScore [7], we use diffusion models in a less compressed domain, which are more expressive generative models than VAEs.
We demonstrate that the proposed speech quality estimation correlates well with traditional intrusive metrics such as POLQA, SI-SDR and SNR. In particular, our method has a higher correlation to these metrics compared to SpeechLMScore on the traditional VoiceBank-DEMAND noisy speech benchmark. Furthermore, we show that, in constrast to SpeechLMScore and VQScore, our method rates utterances processed by speech enhancement baselines in a similar fashion as intrusive and supervised DNN-based non-intrusive metrics.
II Score-based Likelihood Estimation
Score-based generative models [11] are continuous-time diffusion models relying on stochastic differential equations. Such models can learn complex, high-dimensional data distributions such as human speech [13], natural images [10, 11] or music [14]. Score-based models can be considered as iterative Gaussian denoisers. At training time, a so-called forward diffusion process maps the target data distribution to a tractable Gaussian distribution by gradually adding Gaussian-distributed noise. New data is then generated following the reverse diffusion process, which iteratively denoises an initial Gaussian sample until a sample belonging to the target distribution emerges.
Song et al. [11, App. D] show that every stochastic diffusion process has a corresponding deterministic process described by an ODE, whose trajectories share the same marginals as the original diffusion process. This specific ODE is named the probability flow ODE, and it continuously increases (forward in time) or decreases (backward) the level of noise in the data. Karras et al. [15] formulate the probability flow ODE as
where is a noise schedule defining the level of noise at time and is its derivative with respect to . is the score function, a vector field pointing in the direction of higher density of data:
Here, is a denoiser function, implemented as a neural network . In order to stabilize and facilitate the training of the model in spite of varying levels of noise, a series of -dependent scaling operations , , and are used, preconditioning inputs and outputs of the network to have unit variance, as well as a skip connection to avoid amplifying errors in :
When as suggested in Karras et al. [15], Equation 1 can be written as
where we have defined the drift that is central to the calculations in Equations 5, 6 and 8.
Song et al.[11] leverage the probability-flow ODE associated with their stochastic differential equation (SDE) to compute the log-likelihood of the input data. This is similar to the density estimation procedure from neural ODEs, leveraging the instantaneous change of variables formula [16]:
Following Grathwohl et al. [17], by integrating Equation 5 for the ODE in Equation 4, we get the following expression for the log density of the data :
Additionally, the authors show that the trace of the Jacobian matrix can be efficiently computed using the Hutchinson estimator [18]:
where the distribution must satisfy and . This avoids the computation of a separate derivative for each element of the diagonal of the Jacobian matrix.
The following coupled initial value problem is then solved:
with initial value . Both equations are solved simultaneously using the same solver. The second vector entry of the solution then corresponds to the log-likelihood estimate. The vector-Jacobian product can be evaluated at roughly the same cost as a computation of by performing reverse-mode automatic differentiation, having already computed the forward pass when solving for . After solving the system of equations, the value of is obtained via Equation 6 by plugging in the solved value of in the computation of , considering that it follows a multivariate Gaussian distribution.
EARS-WHAM | VoiceBank-DEMAND | |||||||
POLQA | SI-SDR | POLQA | SI-SDR | |||||
Measure | PCC | SRCC | PCC | SRCC | PCC | SRCC | PCC | SRCC |
NISQA [6] | 0.797 | 0.816 | 0.689 | 0.712 | 0.897 | 0.896 | 0.608 | 0.591 |
DNSMOS OVRL [5] | 0.667 | 0.711 | 0.625 | 0.650 | 0.776 | 0.828 | 0.542 | 0.569 |
VQScore [9] | 0.723 | 0.755 | 0.804 | 0.821 | 0.837 | 0.841 | 0.539 | 0.537 |
()SpeechLMScore [8] | 0.761 | 0.779 | 0.733 | 0.760 | 0.702 | 0.681 | 0.471 | 0.428 |
Diffusion Log-likelihood | 0.640 | 0.667 | 0.617 | 0.633 | 0.831 | 0.835 | 0.489 | 0.498 |
III Implementation Details
The framework used in this work to train the diffusion model is the denoising score matching formulation proposed by Karras et al. [15]. Our neural network follows the ADM architecture [19], with the architectural and training improvements subsequently proposed in [20], in particular the magnitude preserving layers. We reduce the model size by using only 3 resolutions and employing one residual block per resolution, resulting in a model with 49M parameters. We set the post-hoc expoential moving average (EMA) length to . For estimation of the trace, we sample from the Rademacher distribution. After experimenting and finding no significant differences with respect to the number of samples, we compute the trace using only one noise vector.
We train the model on the EARS dataset [21], following the train–validation–test split performed for the EARS-WHAM set proposed in the same work [21]. This training set consists of approximately 100 hours of clean, anechoic English speech data with different speaking styles. Here, the audio is downsampled to 16kHz and transformed into mel spectrograms, with 80 mel bands and Hann windows of length 64 ms with 75% overlap. The dynamic range of the spectrogram is compressed using the logarithm, and the values are then scaled to match mean 0 and standard deviation of 0.5, using the statistics computed on the training set. During training, we sample segments of 4 seconds in length, with random starting indices. The segments are sampled from the dataset and train for approximately 37M samples, with batch size 128. We perform evaluations on the EARS-WHAM test set at 16kHz. The test set contains 6 speakers and input signal-to-noise ratios randomly sampled in a range of dB. We additionally report results on the VoiceBank-DEMAND (VB-DMD) test set [22] also downsampled to 16kHz, as a mismatched condition for the model. The test set contains two speakers and noise at at 2.5, 7.5, 12.5 and 17.5dB SNR.
IV Results and Discussion
In Table I, we show correlations with intrusive metrics on the noisy data from the EARS-WHAM and VB-DMD test sets, compared to the other non-intrusive baselines. We report the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC), quantifying linear and monotonic relationships, respectively. We can see that while the correlation of the diffusion-based log-likelihood is generally below that of the other metrics on EARS-WHAM, it performs similarly to VQScore and better than SpeechLMScore on VoiceBank-DEMAND. Figure 1 shows some of these correlations in further detail, with density plots of the diffusion-based log-density over SI-SDR, POLQA and input SNR, alongside a regression line. The plots confirm the correlation, showing a concentration of the samples alongside the diagonal.
Figure 2 shows how the distribution of the log-densities for clean EARS-WHAM test data compare to that of the noisy. The values for clean data have a large standard deviation, which hints that the clean speech in the test set has variations which are not completely modeled by the neural network. Nevertheless, the two distributions are clearly separated, as one would wish for in a quality metric.
One important aspect to analyze is how a measure handles corruptions from an enhancement model, which can have different behaviors depending on the training paradigm [25, 26, 27]. In Table II we show such a comparative evaluation on the EARS-WHAM set, comparing the generative method SGMSE+ [24, 21] against the predictive method Demucs [23], both trained on EARS-WHAM data. To paint a complete picture of the models’ performances, we evaluate them with a set of non-intrusive and intrusive metrics [28]. In the non-intrusive group of metrics, we report VQScore and SpeechLMScore (in perplexity, so lower is better), as well as metrics trained in a supervised fashion, with MOS labels. In the intrusive subset, we employ PESQ [2] and POLQA [3], as well as SI-SDR [1] and phoneme similarity, to spot phonetic confusions typically introduced by generative models [24]. To make it compatible with the metrics, we downsample the enhanced files to 16kHz. Here we can see that, along with DNSMOS and NISQA, log-likelihood favors the SGMSE+ over Demucs, whereas VQScore [9] and SpeechLMScore [8] show a only a tiny difference between these two evaluated enhancement methods, and even produce values slightly in favor of Demucs. The log-likelihood is therefore the only non-intrusive method trained without access to paired MOS data that correctly reflects the clear human listener preference for the method SGMSE+ over Demucs reported in [21]. Furthermore, the log-likelihood also shows better alignment with POLQA and PESQ scores.
V Conclusion
We proposed a speech quality estimator based on unconditional score-based diffusion models trained on clean speech only. Using the natural likelihood computation abilities of score-based models, the proposed estimator can estimate the quality of speech utterances without having any access to paired data. The resulting measure is non-intrusive and yet correlates well with intrusive metrics on noisy speech benchmarks. When evaluating utterances processed by speech enhancement baselines, our method is the only unsupervised non-intrusive metric that properly correlates with other intrusive metrics, as well as supervised DNN-based non-intrusive models.
VI Acknowledgement
We would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.